Internal Rate of Return Method—Two Projects

Cousin's Salted Snack Company is considering two possible investments: a delivery truck or a bagging machine. The delivery truck would cost $90,790.20 and could be used to deliver an additional 47,000 bags of taquitos chips per year. Each bag of chips can be sold for a contribution margin of $0.58. The delivery truck operating expenses, excluding depreciation, are $0.79 per mile for 16,000 miles per year. The bagging machine would replace an old bagging machine, and its net investment cost would be $60,018.75. The new machine would require three fewer hours of direct labor per day. Direct labor is $15 per hour. There are 250 operating days in the year. Both the truck and the bagging machine are estimated to have eight-year lives. The minimum rate of return is 9%. However, Cousin's has funds to invest in only one of the projects.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Compute the internal rate of return for each investment. Use the above table of present value of an annuity of $1. Enter the present value factor to three decimal places.

Delivery Truck Bagging Machine
Present value factor
Internal rate of return % %

b. Which of the following should be included in a memo to management regarding the recommendation about the two projects?
Answer #1
Net annual cash flows:
Delivery Truck 14620 =(47000*0.58)-(16000*0.79)
Bagging Machine 11250 =250*3*15
a
Delivery Truck Bagging Machine
Present value factor 6.210 5.335
Internal rate of return 6% 10%
b
Bagging Machine is recommended as it's Internal rate of return is higher
Workings:
Delivery Truck Bagging Machine
Present value factor =90790.2/14620 =60018.75/11250